a roofer spends 12 hours to shingle a roof alone. The roofer quits after 3 hours and the helper spends another 12 hours to finish. how long would it take the helper alone to do the whole job?
Work Attempted:
12/3=(12+x)/15
I have tried a few different ways and keep coming back to an answer of 60 hours; I am not comfortable with this answer.

Answer by nerdybill(7384)

a roofer spends 12 hours to shingle a roof alone. The roofer quits after 3 hours and the helper spends another 12 hours to finish. how long would it take the helper alone to do the whole job?
.
Let x = time (hours) it takes for helper working alone
.
roofer:
1 roof per 12 hours
.
helper
1 roof per x hours
.
3(1/12) + 12(1/x) = 1
multiplying both sides by 12x:
3(x) + 12(12) = 12x
3x + 144 = 12x
144 = 9x
16 hours = x